OpenStudy (anonymous):

Ethylene glycol (antifreeze) has a density of 1.11 g/cm^3. 1/What is the mass in grams of 385mL of this liquid? 2/What is the volume in liters of 3.43 kg of this liquid? Please it is due in one hour!

OpenStudy (anonymous):

M = D*V cm^3=mL so (1.11 g/cm^3)* 385 mL ^ but we cant have cm^3 so we replace it by cm^3=mL (1.11 g/(mL))* 385 mL =427.35 g

OpenStudy (anonymous):

2 What is the volume in liters of 3.43 kg of this liquid? density of 1.11 g/cm^3. M/D=V Plug in info (3.43 kg) / (1.11 g/cm^3)=V ^ not right units, suppose to be in g so move over 3 to right and you get it in g (3430 g) / (1.11 g/cm^3)=3090.090090*cm^3 put you want it in litters 3090.090090*cm^3 ^ replace by mL 3090.090090 mL ^ replace by 1 L= 1000 mL -> 1/1000 L=mL 3.090090090 move decimal 3 places to left

OpenStudy (anonymous):

so the final answer for the second part is 3.090090090?

OpenStudy (anonymous):

rougly 3 liters depends on sigfigs 3.09 L

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

de nada